# Distribution Matching Distillation Meets Reinforcement Learning

### v3

@@ 4.1 Experimental Setup
We provide a brief overview of our training and evaluation schemes here.
For training, we adopt denoising-based model (SDXL-Base [sdxl]) and flow-based models (SD3-Medium [sd3], SD3.5-Large [sd35]) for distillation, using prompts from t2i-2M [t2i-2M]. Additionally, we use ReFL [refl] with DFN-CLIP [dfnclip] and HPSv2.1 [hpsv2] as reward models by default. As for evaluation, we validate the effect of DMDR through extensive experiments. First, we use prompts sampled from a recent public dataset ShareGPT-4o-Image [sharegpt4o-image] to generate images and follow previous work [dmdx, tdm] that report CLIP Score [clipscore], Aesthetic Score [laion5b], Pick Score [pickscore], and Human Preference (HP) Score [hpsv2]. Next, in order to obtain a more comprehensive evaluation, we also compare our distilled models with their teachers on two popular benchmarks, DPG_Bench [DPG] and GenEval [geneval]. Finally, we conduct both qualitative and quantitative ablation experiments to further verify our method’s effectiveness.
Table 1: System-level comparison against state-of-the-art methods. denotes our reproduced results. Best performance is marked in Bold. Image-Free denotes whether the training requires external image data or data generated by the multi-step teacher.
@@
Base-Model (CFG = 7.0) 25 50 34.7588 5.6480 22.1085 27.1477 -
LCM [lcm] (CFG = 8.0) 1 2 28.4664 5.1026 20.0603 17.6837 $\times$
Lightning [sdxl-lightning] 1 1 32.0283 5.6761 21.4868 26.3615 $\times$
DMD2 [dmd2] 1 1 34.3046 5.6238 21.7293 26.5986 $\times$
DMDR (ours) 1 1 35.4835 6.0483 22.5424 31.1442 $\checkmark$
DMD2 [dmd2] 4 4 34.5169 5.7043 22.1546 28.5655 $\times$
DMDR (ours) 4 4 35.2940 5.9857 22.6268 32.8678 $\checkmark$
@@
Base-Model (CFG = 7.0) 25 50 34.9025 5.5942 22.1801 28.4021 -
Hyper-SD [hypersd] (CFG = 5.0) 4 8 32.0234 5.2489 20.2831 22.4544 $\times$
DMD2 [dmd2] 4 4 33.9421 5.6137 21.7688 27.3675 $\times$
Flash [flashdiff] 4 4 34.2634 5.6702 21.5921 26.6542 $\times$
TDM [tdm] 4 4 34.0301 5.6250 22.0010 27.7522 $\checkmark$
DMDR (ours) 4 4 34.9542 5.8462 22.3578 31.8979 $\checkmark$
@@
Base-Model (CFG = 3.5) 25 50 35.5509 5.7014 22.4856 28.8135 -
LADD [ladd] 4 4 35.0480 5.4514 22.2451 27.8470 $\times$
DMDR (ours) 4 4 35.8647 6.0284 22.8859 32.4724 $\checkmark$
Table 2: Quantitative comparison on DPG_Bench [DPG] of our four-step model against its multi-step teacher.
@@
Ours 85.30 90.46 90.50 90.66 87.40 90.18
4.2 Main Results
We begin by assessing DMDR’s text-to-image generation performance using prompts sampled from ShareGPT-4o-Image [sharegpt4o-image], ensuring that there is no overlap with the prompts utilized during training. Table 4.1 presents a summary of the distillation results in comparison to other methods. Models distilled through our approach exhibit state-of-the-art (SOTA) capabilities in terms of prompt coherence and the generation of high-quality, aesthetically pleasing images. Notably, our method demonstrates effectiveness across various architectures (e.g., UNet, Transformer), model sizes (e.g., 2B, 8B), and paradigms (e.g., flow-based, denoising-based), highlighting the broad applicability and universality of the proposed approach. Additionally, we provide qualitative comparisons in Figure 5 and Appendix B, where our method produces images characterized by superior quality and enhanced prompt coherence.
To further verify our method’s effectiveness and provide a more comprehensive comparison between the few-step distillation model and the multi-step teacher model, we conduct the comparison on two commonly used benchmarks [DPG, geneval], Table 2 shows our four-step model consistently outperforms its multi-step teacher on DPG_Bench across all base models, lifting the overall score from 74.65 to 76.44 for SDXL-Base, 84.08 to 84.96 for SD3-Medium, and 84.12 to 85.30 for SD3.5-Large. Meanwhile, on GenEval (Table 3) the same trend holds: overall scores improve for SDXL (0.55 → 0.56), SD3 (0.62 → 0.64), and SD3.5 (0.71 → 0.72). These results indicate that DMDR has successfully freed the student model from the constraints of the multi-step teacher model and stimulated its capabilities during the distillation process.
Table 3: Quantitative comparison on GenEval [geneval] of our four-step model against its multi-step teacher.
Model Overall Single Two Count. Colors Pos. Attr.
SDXL-Base
Teacher 0.55 0.98 0.74 0.39 0.85 0.15 0.23
Ours 0.56 0.99 0.76 0.42 0.84 0.11 0.24
SD3-Medium
Teacher 0.62 0.98 0.74 0.63 0.67 0.34 0.36
Ours 0.64 0.99 0.84 0.54 0.81 0.25 0.44
SD3.5-Large
Teacher 0.71 0.98 0.89 0.73 0.83 0.34 0.47
Ours 0.72 0.99 0.93 0.68 0.80 0.30 0.59
Figure 7: Ablation examples of two dynamic training strategies. Better to zoom in to check the effect.
@@ 4.3 Ablation Study
We investigate our method’s properties below. We use SD3-Medium with $512\times 512$ resolution by default for a fast ablation.
Effect of Dynamic Cold Start Stage. As shown in Table 4.3, our dynamic cold start strategies significantly enhance the initial training phase. Firstly, introducing DynaDG leads to a notable improvement. By dynamically adjusting the $P_{real}$ to create more overlap with the nascent $P_{fake}$ distribution, DynaDG provides more reliable gradient signals. Quantitatively, this translates to a CLIP Score increase from 31.56 to 32.79 (+1.23) and an HP Score increase from 27.01 to 27.57 (+0.56) at 1000 iterations. Similar gains are observed at 2000 iterations. Further incorporating DynaRS, which initially biases renoise sampling towards higher noise levels, guiding the generator to learn global structures before progressively shifting to finer details, further boosts performance. To highlight the importance of the dynamic aspect of our strategies, we evaluate a variant denoted as -Dynamic. This variant removes the dynamic scaling of LoRA in DynaDG (e.g., using a fixed LoRA scale) and employs biased renoise sampling in the whole training process. While it still outperforms vanilla_DMD, it consistently underperforms the full dynamic approaches. This clearly indicates that the adaptive, progressive nature of our DynaDG and DynaRS mechanisms is crucial for maximizing performance during the cold start phase. Qualitative results presented in Figure 7 further corroborate these findings, showing a faster global structure building when our dynamic cold start strategies are employed.
Table 4: Ablation on our dynamic cold start stage.
Method 1000 iter 2000 iter
CLIP Score HP Score CLIP Score HP Score
vanilla_DMD 31.5608 27.0051 32.5492 28.4541
   +DynaDG 32.7905 27.5723 33.5045 29.0363
   +DynaRS 33.1657 27.8276 33.6432 29.1157
   -Dynamic 32.6562 27.2641 33.0245 28.8459
Effect of DMDR with Different RL Algorithms. We choose three representative RL methods [refl, diffusiondpo, flowgrpo] to illustrate the efficacy of combining distillation with various RL techniques in DMDR, the results are provided in Table 4.3. Firstly, we observe that the initial model state after our dynamic cold start stage already provides a good baseline performance. Training with “w/ only Distill.”, as we illustrated above, this would be inherently capped by the teacher’s distribution, resulting in a very minor improvement over the initial state. Next, when considering introducing different RL algorithms (the regularization originally used in these methods has also been added), although these show improvements in some metrics, they do not consistently surpass the “only Distill” baseline across all scores. Moreover, the reward hacking occurs (we will analyze it specifically in the next part). Crucially, the rows corresponding to “w/ Distill. + RL” consistently demonstrates superior performance across all evaluated metrics against the counterparts. Figure 8 also shows superior visualization results when DMD and RL are conducted together. These validate our core insight that RL unlocks the performance potential of DMD, allowing the student model to surpass the teacher’s inherent limitations by optimizing for external reward signals. Simultaneously, DMD effectively regularizes the RL process, preventing reward hacking and ensuring that the student’s distribution remains robust and comprehensive, as evidenced by the consistent improvements across diverse metrics with different RL algorithms.
Table 5: Ablation on RL and DMD. We init the model after cold start for 2K iteration then continue training in different ways.
Method CLIP Aesthetic Pick HP
Score Score Score Score
init 33.6432 5.6124 21.0489 29.1157
w/ only Distill. 33.6738 5.6248 21.6376 29.1389
w/ only RL (ReFL) 33.1897 5.8841 22.3008 31.2714
w/ Distill. + RL (ReFL) 34.6249 6.1813 22.7578 32.8979
w/ only RL (DPO) 33.9554 5.7893 21.7643 29.5819
w/ Distill. + RL (DPO) 33.9632 5.9710 21.9865 30.5994
w/ only RL (GRPO) 33.9095 5.7876 21.7956 29.4429
w/ Distill. + RL (GRPO) 34.0055 5.8256 22.0120 30.6248
Figure 8: Ablation examples of different RL algorithms with DMD. Better to zoom in to check the effect.
@@
Furthermore, an interesting observation is that using the ReFL style RL method achieves overall better performance, and our analysis is that: traditional ReFL for multi-step models ignores earlier sampling steps and only trains the last few steps before the output image [refl] because the large computational graph generated by the gradient of the reward model and the multi-step denoising process would result in a huge memory consumption. Subsequent works [srpo, deepreward] mainly focus on saving forward noise trajectory and sampling a subset of steps to optimize to remedy this dilemma, although this can allow the gradient to return to the early denoising step, it is also prone to error accumulation. On the contrary, in our few-step model, the model is trained to predict a clean image at each step’s state [dmd2], which enables the gradient of ReFL to directly return to the initial state, thus obtaining an effective optimization.
Figure 9: Reward curves comparison during the RL stage. Better to zoom in to check the difference.
Effect of DMDR to Mitigate Reward Hacking. We follow RewardDance [rewarddance] to present the reward curves during training illustrated in Figure 9. It is evident that employing DMD loss as a form of regularization (right) demonstrates significantly greater variance throughout the training process compared to the traditional approach (left). This increased diversity in reward signals suggests that models trained using our method exhibit enhanced robustness against reward hacking. This assertion is further supported by Figure 8, which shows that in the absence of DMD loss, utilizing ReFL results in a rasterized appearance of images, indicating the occurrence of reward hacking.
5 Limitation and Future Work
We discuss some limitations we realized during the development of DMDR. We hope these will serve as a catalyst for innovative ideas and advancements in future research.
Quality vs. Diversity. One notable thing is the trade-off between quality and diversity. Although models trained from our method show superior quality and prompt coherence, the diversity is reduced like shown in Table 6. We believe that this is a common feature of score distillation like DMD and RL, that is, sacrificing uncertainty in sampling in exchange for stable and desired results, not a limitation unique to our approach. The reduced diversity caused by DMD, we believe, will also have some impact on some online RL methods that require a group of samples for ranking then optimization. This is also reflected in the slightly poorer improvement of GRPO and DPO compared to ReFL in the Table 4.3. Further investigation into finding the optimal balance between quality and diversity is left for future work.
Table 6: Diversity evaluation on ShareGPT-4o-Image [sharegpt4o-image].
DMD DMD2 Ours Teacher
[…]
Scaling of Reward Model. Another notable thing is the scaling of the reward model in DMDR. Our experimental results demonstrate that our method enables the few-step model to be able to surpass its teacher model. However, as indicated in the Table 2 and 3, certain metrics exhibit a decline after the distillation. We attribute this degradation primarily to inherent limitations and potential biases present in the open-source reward model we employed. Notably, RewardDance [rewarddance] has shown that scaling up the reward model yields consistent improvements for RL in multi-step diffusion models. We posit that similar benefits would extend to few-step models with our method, but require more computation resources, which we leave for future work.
6 Conclusion
This work introduces DMDR, which combines Distribution Matching Distillation (DMD) with reinforcement learning (RL) in one framework. We show that DMD and RL can be trained simultaneously, with RL enabling the student model to surpass the teacher and DMD loss regularizing RL to prevent reward hacking. We also propose dynamic training strategies (DynaDG and DynaRS) to enhance the initial distillation process. Experiments demonstrate state-of-the-art performance among few-step methods and overall superior performance against multi-step teachers.
Appendix
Appendix A More Related Works

### v4

@@ 3.1 Experimental Setup
For training, we adopt denoising-based model (SDXL-Base [sdxl]) and flow-based models (SD3-Medium [sd3], SD3.5-Large [sd35]) for distillation, using prompts from t2i-2M. Additionally, we use ReFL [refl] with DFN-CLIP [dfnclip] and HPSv2.1 [hpsv2] as reward models by default. As for evaluation, we validate the effect of DMDR through extensive experiments. First, we use prompts sampled from a recent public dataset ShareGPT-4o-Image [sharegpt4o-image] to generate images and follow previous work [dmdx, tdm] that report CLIP Score [clipscore], Aesthetic Score [laion5b], Pick Score [pickscore], and Human Preference (HP) Score [hpsv2]. Next, in order to obtain a more comprehensive evaluation, we also compare our distilled models with their teachers on two popular benchmarks, DPG_Bench [dpg] and GenEval [geneval]. Noted that there are some differences from the settings we made in the Method section. We train 1.5K steps in stage1 and stage 2, respectively, to achieve better performance.
Table 2: System-level comparison against state-of-the-art methods. denotes our reproduced results based on the official code. Best performance is marked in Bold.
@@
Base-Model 50 - 34.7588 5.6480 22.1085 27.1477
ReFL [refl] 50 RL 35.2107 5.9045 22.8076 33.1874
LCM [lcm] 2 Distill 28.4664 5.1026 20.0603 17.6837
Lightning [sdxl-lightning] 1 Distill 32.0283 5.6761 21.4868 26.3615
DMD2 [dmd2] 1 Distill 34.3046 5.6238 21.7293 26.5986
DMDR (ours) 1 Distill & RL 35.6241 6.1184 22.8498 32.0065
DMD2 [dmd2] 4 Distill 34.5169 5.7043 22.1546 28.5655
DMD2-PSO [pso] 4 Distill then RL 34.0128 5.8032 22.2644 29.8732
DMDR (ours) 4 Distill & RL 35.4940 6.0324 22.7122 32.9832
@@
Base-Model 50 - 34.9025 5.5942 22.1801 28.4021
ReFL [refl] 50 RL 35.1851 5.7821 22.0064 32.0745
DMD2 [dmd2] 4 Distill 33.9421 5.6137 21.7688 27.3675
Flash [flashdiff] 4 Distill 34.2634 5.6702 21.5921 26.6542
TDM [tdm] 4 Distill 34.0301 5.6250 22.0010 27.7522
Hyper-SD [hypersd] 8 Distill then RL 32.0234 5.2489 20.2831 22.4544
DMDR (ours) 4 Distill & RL 35.0142 5.8876 22.4892 32.1145
@@
Base-Model 50 - 35.5509 5.7014 22.4856 28.8135
LADD [ladd] 4 Distill then RL 35.0480 5.4514 22.2451 27.8470
DMDR (ours) 4 Distill & RL 35.9757 6.1541 22.9072 32.7368
Method for comparison. We compare our method against several categories of existing work, including foundational multi-step base models [sdxl, sd3, sd35], RL only approaches [refl], distillation only methods [lcm, sdxl-lightning, dmd2, flashdiff, tdm], distillation then RL frameworks [hypersd, pso]. We do not compare with methods like Flash-DMD [flash-dmd] and Diff-Instruct++ [diffinstruct++] that do not provide model weights or open source training code because we cannot guarantee a fair and reproducible evaluation.
3.2 Main Results
We begin by assessing DMDR’s text-to-image generation performance using prompts sampled from ShareGPT-4o-Image [sharegpt4o-image], ensuring that there is no overlap with the prompts utilized during training. Table 3.1 presents a summary of the distillation results in comparison to other methods. Models distilled through our approach exhibit state-of-the-art (SOTA) capabilities in terms of prompt coherence and the generation of high-quality, aesthetically pleasing images. Notably, our method demonstrates effectiveness across various architectures (e.g., UNet, Transformer), model sizes (e.g., 2B, 8B), and paradigms (e.g., flow-based, denoising-based), highlighting the broad applicability and universality of the proposed approach. Additionally, we provide qualitative comparisons in Figure 7, where our method produces images characterized by superior quality and enhanced prompt coherence.
Table 3: GenEval Comparison: 4-step (ours) $vs.$ its multi-step teacher.
Model Overall Single Two Count. Colors Pos. Attr.
SDXL-Base
Teacher 0.55 0.98 0.74 0.39 0.85 0.15 0.23
Ours 0.57 0.99 0.76 0.44 0.85 0.12 0.25
SD3-Medium
Teacher 0.62 0.98 0.74 0.63 0.67 0.34 0.36
Ours 0.65 0.99 0.84 0.57 0.81 0.27 0.45
SD3.5-Large
Teacher 0.71 0.98 0.89 0.73 0.83 0.34 0.47
Ours 0.72 0.99 0.93 0.69 0.80 0.31 0.59
Table 4: DPG_Bench Comparison: 4-step (ours) $vs.$ its multi-step teacher.
@@ 3.2 Main Results
Ours 85.33 90.47 90.53 90.68 87.44 90.20
Figure 7: Visual comparison between the teachers, selected competing methods [dmd2, hypersd, ladd], and ours. All images are generated using identical noise. Our model produces images with superior quality and prompt coherence.
To further verify our method’s effectiveness and provide a more comprehensive comparison between the few-step distillation model and the multi-step teacher model, we conduct the comparison on two commonly used benchmarks [dpg, geneval], Although we do not perform RL optimization on the metric of these two benchmarks, table 4 and  4 show that the overall score of our four-step model consistently outperforms its multi-step teacher on DPG_Bench and GenEvalacross all base models. These results indicate that DMDR has successfully freed the student model from the constraints of the multi-step teacher model and stimulated its capabilities during the distillation process.
Table 5: Ablation on different components. Default settings are marked in purple.
(a) Ablation on different reward models.
Method CLIP Aesthetic Pick HP DPG
Score Score Score Score overall
Pick [pickscore] 34.3403 5.9021 22.9674 31.7665 84.86
AE [laion5b] 34.1748 6.1024 22.3858 31.6620 84.61
CLIP [clipscore] 35.2087 5.5832 22.0644 28.0507 85.10
HPS [hpsv2] 34.4088 5.9021 22.5108 32.9065 85.04
CLIP + HPS 35.0142 5.8876 22.4892 32.1145 85.30
(b) Ablation on RL algorithm in stage 2.
Method CLIP Aesthetic Pick HP DPG
Score Score Score Score overall
stage 1 init 33.3648 5.5744 21.0070 28.8371 82.4371
w/ RL (ReFL) 35.0142 5.8876 22.4892 32.1145 85.30
w/ RL (DPO) 34.0234 5.8340 21.9844 30.8352 85.00
w/ RL (GRPO) 34.6675 5.9324 22.3248 31.4474 85.08
(c) Ablation on Value of $\lambda_{\text{rl}}$.
Value CLIP Aesthetic Pick HP DPG
Score Score Score Score overall
0.1 34.7422 5.8508 22.3064 31.6565 85.24
0.5 35.0142 5.8876 22.4892 32.1145 85.30
1.0 35.0087 5.84430 22.4709 32.8982 85.04
2.0 35.0322 5.7806 22.1098 32.7683 82.48
@@ 3.3 Ablation Study
In this section, we conduct a systematic investigation into the key components of DMDR. Unless otherwise specified, all experiments use the SD3-M base model with a 4-step sampling configuration.
Versatility across reward models. A key advantage of ourusing a DINOv2-based reward model to maximize classification accuracyDMDR is the ability to steer the distillation process toward diverse human preferences. As shown in Table 5(a), we evaluate the framework using different reward models (RMs) targeting specific attributes. The results provide a clear insight: optimizing with a specific RM leads to a significant performance gain in its corresponding metric, confirming that DMDR effectively injects the desired preference signal into the few-step manifold. Furthermore, by integrating multiple rewards (CLIP + HPS), the model achieves a more balanced and superior generative performance, similar to the findings in DanceGRPO [dancegrpo]. This flexibility allows users to customize the distillation ”flavor” according to specific downstream tasks.
@@
Compatibility with RL paradigms. To verify the algorithmic universality of our unified framework, we compare three distinct RL strategies in the Stage 2 joint optimization phase: ReFL [refl], DPO [diffusiondpo, pso], and GRPO [dancegrpo, flowgrpo]. As shown in Table 5(b), all three algorithms yield substantial improvements over the ”Stage 1 init” baseline across all evaluation dimensions. Furthermore, an interesting observation is that using the ReFL in DMDR achieves overall better performance, and our analysis is that: traditional ReFL for multi-step models ignores earlier sampling steps and only trains the last few steps before the output image [refl] because the large computational graph generated by the multi-step denoising process would result in a huge memory consumption. Subsequent solutions [srpo, deepreward] mainly focus on saving forward noise trajectory and sampling a subset of steps to optimize, although this can allow the gradient to return to the early denoising step, it is also prone to error accumulation. On the contrary, in our few-step model, the model is trained to predict a clean image at each step’s state [dmd2], which enables the gradient of ReFL to directly return to the initial state, thus obtaining an effective optimization.
The balancing Role of DMD Regularization. The synergy between distillation and RL is governed by the balancing coefficient $\lambda_{\text{rl}}$ in our joint objective. Table 5(c) illustrates the trade-off between reward maximization and distributional stability. When $\lambda_{\text{rl}}$ is set too low; the model remains constrained by the teacher’s manifold, limiting the potential for preference alignment. Conversely, as it increases beyond an optimal threshold, we observe a noticeable decline in a general capacity (reflected by the DPG overall score), despite high reward scores. This trend provides a crucial insight: the DMD loss serves as a vital structural regularizer. It prevents the model from ”reward hacking”(We also showed in Figure 3. The optimal balance ensures that the model pushes the boundaries of quality while staying anchored to the teacher’s robust generative priors.
4 Limitation and Discussion
Quality vs. Diversity. To further analyze the behavior of DMDR, we conduct class-conditional experiments on ImageNet $\times256$ using the SiT [sit] backbone with representation alignment pretraining ckpt [repa, sra] as initialization for 1-step generation. As shown in Table 6, a fundamental trade-off between generative quality and diversity emerges. Our results reveal that both Classifier-Free Guidance (CFG) (this also observed in Decoupled-DMD [decoupled-dmd]) and Reinforcement Learning (here we use a DINOv2-based [dinov2] reward model to maximize classification accuracy) act as ”density sharpeners” during the distillation process. This phenomenon suggests that DMDR effectively ”sacrifices” the teacher’s distributional uncertainty in exchange for the stability and quality required for few-step sampling. We believe this reduction in diversity is not a limitation unique to our approach, but a characteristic of score-based distillation and preference alignment, which steer the model toward high-reward, high-confidence regions. This trade-off is often desirable in practical few-step applications where sample excellence and prompt adherence are prioritized over exhaustive distribution coverage. But at the same time, further investigation into finding the optimal balance between quality and diversity is left for future work.
Table 6: 1-step SiT DMDR Results on ImageNet $256\times 256$. We report the performance with different CFG scales and RL optimization. Metrics include FID [fid], Recall [pre-rec], Inception Score (IS) [is], and Precision [pre-rec].
CFG RL FID $\downarrow$ Recall $\uparrow$ Inception Score $\uparrow$ Precision $\uparrow$
No No 2.13 0.64 232.15 0.77
1.5 No 5.20 0.48 387.24 0.88
4 No 15.70 0.11 453.65 0.86
No Yes 6.95 0.40 416.01 0.90
1.5 Yes 13.43 0.24 494.43 0.93
5 Related Work
Distribution matching distillation. Distribution Matching Distillation [dmd] (DMD) is the first work to successfully apply the principle score-based distillation [score-base-diff, prolificdreamer, dreamfusion] to large-scale text-to-image models. Intuitively, it strives to ensure that any sample realized by the student at a given noise level occurs with exactly the same probability as it would under the teacher’s distribution, thereby preserving the multi-step generative priors in the few-step model. From then on, a lot of follow-up work emerged [dmd2, dmdx, tdm, sd3.5flash, f-distill, flash-dmd, twinflow], for example, DMD2 [dmd2] employs a discriminator to align the student model distribution with a specific target distribution like GAN [gan] does; TDM [tdm] incorporates DMD loss in the sampling process of the student model for better alignment; ff–distill [f-distill] replaces the original reverse Kullback–Leibler (KL) divergence to the proposed $f$-divergence for covering different divergences with different properties. Our work also starts with DMD, but we don’t focus on how to better “imitate” the teacher. Instead, we aim to incorporate reinforcement learning in the distillation process to enable a more controllable and preference-aligned distillation.
Reinforce learning for diffusion models. Reinforce learning helps to align diffusion models to human preferences by training a reward model and using it to guide generation [ddpo, flowgrpo, refl, srpo, diffusiondpo, dancegrpo, diffusionnft]. There are many algorithms to conduct RL, for example, DDPO [ddpo] adapts PPO via image log-likelihoods; ReFL [refl] bypasses likelihoods by optimizing outputs with frozen-reward gradients; Diffusion-DPO [diffusiondpo] adapts DPO to diffusion for paired human preference data; and recent GRPO extensions [flowgrpo, dancegrpo] use GRPO to diffusion models by coupling the training loss with SDE samplers. Although significant progress has been made in RL for diffusion models, most of the work has focused on multi-step models. Here, we find that these algorithms are also adapted to the few-step model when carried out in conjunction with distribution matching distillation.